# A new method for optical steel rope non-destructive damage detection

### v2

@@ VI Conclusion
This paper proposes a dataset for the segmentation and detection of the steel ropes used in amusement facility and elevator related scenarios. A new segmentation model named RGBD-UNet is proposed, which integrates the CMA module into the U-Net to effectively fuse color and depth information. We also propose a new detection model named VovNetV3.5, which combines the DBB module into VovNetV2. And a series of experiments are conducted to prove the advantages of our models.
In future research endeavors, it might be beneficial to consider incorporating CMA modules and depth information into other U-structure models like U2net or BasNet[36]. Currently, due to time and human resources limitations, the classification of steel ropes is restricted to two broad categories: normal and abnormal. However, there is potential for future studies to delve into a more granular classification of abnormal steel ropes, distinguishing them into subcategories such as twisted, worn, frayed, and rust. Then it could yield more detailed insights into the condition of steel ropes.
Acknowledgment
This research was funded by the Ministry of Science and Technology, the National Key Research and Development Program of China, grant number 2022YFC3005100, and the State Administration for Market Regulation of China, grant number 2022MK211.

### v4

@@ VI Conclusion
In this paper, we propose a dataset for the segmentation and detection of the steel ropes used in amusement facilities and elevator-related scenarios. A new segmentation model named RGBD-UNet is proposed, which integrates the CMA module into the U-Net to effectively fuse color and depth information. We also propose a new detection model named VovNetV3.5, which combines the DBB module into VovNetV2. A series of experiments are conducted to prove the advantages of our models.
In future research endeavors, it might be beneficial to consider incorporating CMA modules and depth information into other U-structure models. Currently, due to time and human resources limitations, the classification of steel ropes is restricted to two broad categories: normal and abnormal. However, there is potential for future studies to delve into a more granular classification of abnormal steel ropes, distinguishing them into subcategories such as twisted, worn, frayed, and rust. Then it could yield more detailed insights into the condition of steel ropes.